# MGSR: 2D/3D Mutual-boosted Gaussian Splatting for High-fidelity Surface Reconstruction under Various Light Conditions

### v1

@@ Abstract
Novel view synthesis (NVS) and surface reconstruction (SR) are essential tasks in 3D Gaussian Splatting (3D-GS). Despite recent progress, these tasks are often addressed independently, with GS-based rendering methods struggling under diverse light conditions and failing to produce accurate surfaces, while GS-based reconstruction methods frequently compromise rendering quality. This raises a central question: must rendering and reconstruction always involve a trade-off? To address this, we propose MGSR, a 2D/3D Mutual-boosted Gaussian splatting for Surface Reconstruction that enhances both rendering quality and 3D reconstruction accuracy. MGSR introduces two branches—one based on 2D-GS and the other on 3D-GS. The 2D-GS branch excels in surface reconstruction, providing precise geometry information to the 3D-GS branch. Leveraging this geometry, the 3D-GS branch employs a geometry-guided illumination decomposition module that captures reflected and transmitted components, enabling realistic rendering under varied light conditions. Using the transmitted component as supervision, the 2D-GS branch also achieves high-fidelity surface reconstruction. Throughout the optimization process, the 2D-GS and 3D-GS branches undergo alternating optimization, providing mutual supervision. Prior to this, each branch completes an independent warm-up phase, with an early stopping strategy implemented to reduce computational costs. We evaluate MGSR on a diverse set of synthetic and real-world datasets, at both object and scene levels, demonstrating strong performance in rendering and surface reconstruction.
@@ 1 Introduction
3D Gaussian Splatting (3D-GS) [1] has recently gained great attention in computer graphics and 3D vision [2]. By representing scenes as collections of 3D Gaussian primitives, 3D-GS offers a more flexible and adaptive representation, enabling accurate and efficient rendering and visualization without relying on neural networks. 3D-GS effectively resolves the issues of low training efficiency and insufficient geometric accuracy in previous NeRF methods. Due to its impressive rendering efficiency and accuracy, 3D-GS has been widely applied in fields such as SLAM [3], dynamic real-world or large-scale scene modeling [4, 5], reconstruction [6, 7], navigation [8], manipulation [9, 10, 11], 3D generation [12, 13], and 3D human simulation [14], etc.
@@
Figure 1: MGSR achieves strong NVS and SR results compared with methods based on 2D-GS [15] and 3D-GS [16]. The input consists of multi-view images captured from various camera positions and angles, under significantly varying light conditions. In some views, the object appears well-lit and clearly visible, while in others, it is poorly lit, resulting in shadows, reflections, and inconsistencies across the images.
@@
Nevertheless, achieving high-fidelity surface reconstruction (SR) and improving the realism of novel view synthesis (NVS) under various light conditions are two main challenges in 3D-GS. Specifically, considering that GS-based rendering methods are affected by different light conditions, previous methods [17, 18] proposed to decompose light to enhance the neural rendering in scenes containing reflective surfaces with the introduction of BRDF (Fig. 1a). However, despite the effectiveness of illumination decomposition in rendering, these methods are time-consuming and still struggle to achieve meaningful mesh extraction due to inherent limitations in 3D-GS. For instance, the centers of Gaussian primitives do not align well with the surface. On the other hand, GS-based SR methods [19, 15, 16] endeavor to constrain the Gaussian primitives close to the surface, inevitably sacrifice rendering quality and are sensitive to light conditions, which may cause surface artifacts on surfaces (Fig. 1b). For example, 2D-GS [15] utilize 2D Gaussian primitives for 3D scene representation, facilitating accurate and view-consistent geometry modeling. However, 2D-GS still cannot effectively model surfaces when ambient lighting changes.
@@
To solve these contradictions, we propose MGSR, a 2D/3D Mutual-boosted Gaussian splatting for Surface Reconstruction that enhances both rendering quality and 3D reconstruction accuracy (Fig. 1c). Specifically, two branches are devised based on 2D-GS and 3D-GS, which are optimized synchronously and mutually enhance each other: the 2D-GS branch provides accurate geometric information, with its limited rendering capabilities improved by 3D-GS branch, while the 3D-GS branch focuses on rendering, with the cost of geometric accuracy compensated by the 2D-GS branch. With geometry information, a geometry-guided illumination decomposition module is devised in the 3D-GS branch to obtain the reflected and transmitted components, achieving realistic rendering under various light conditions. To achieve this, an additional Spherical Harmonic (SH) is introduced to model reflected colors, along with two other parameters: reflected opacity and reflected confidence, to represent the reflected component accurately. Using the transmitted component as supervision, the 2D-GS branch can achieve high-fidelity SR while avoiding the impact of illumination factors on surface estimation. Throughout the mutual-boosted supervision stage, the 2D-GS and 3D-GS branches engage in alternating optimization for mutual supervision. Prior to alternating optimization, the two modules undergo an independent warm-up stage, and an auto-stop strategy is introduced to reduce unnecessary computational burdens. To the best of our knowledge, MGSR is the first GS-based approach that investigates the simultaneous enhancement of rendering and reconstruction, as well as the first mutual-boosted work on GS involving both 2D-GS and 3D-GS. MGSR is thoroughly evaluated across a wide range of synthetic and real-world datasets, as well as object- and scene-level datasets, showcasing its superior performance in both rendering and mesh extraction. In summary, our main contributions are as follows:
@@
  • The 2D-GS branch aims to provide geometry information to enhance the illumination decomposition of the 3D-GS. The decomposed transmitted color will be in turn utilized to supervise the 2D-GS branch for improved surface reconstruction, independent of varying light conditions.
  • Leveraging geometric information from the 2D-GS branch, we have developed a geometry-guided illumination decomposition module to enhance the accuracy of decomposing reflected and transmitted components, achieving more realistic rendering outcomes.
@@ 2 Related work
Recently, several attempts have also been made at lighting estimation and decomposition on 3D-GS. R3DG [17] leverages 3D-GS and NEILF [21] to create a ray-tracing and relighting-capable 3D-GS representation. GShader [18] presents a simplified shading function for reflective surfaces in 3D-GS, where a novel normal estimation is introduced that uses the shortest axis direction of the 3D Gaussian as an approximate normal, eliminating reliance on depth map priors and avoiding the flattening or 2D projection of 3D Gaussians. GS-IR [24] utilizes depth-derivation-based regularization for normal estimation and a baking-based method for modeling indirect lighting, enabling a precise decomposition of material attributes and illumination, thereby significantly improving the photorealism of the rendered outcomes. The prevailing focus in current efforts on lighting estimation and decomposition is predominantly on achieving photorealistic rendering, which consequently hinders the ability to carry out surface mesh extraction effectively.
3D-GS driven surface reconstruction. The misalignment of the centroids (means) of 3D Gaussians with the actual surfaces presents a challenge for accurately reconstructing surfaces. A common strategy is to flatten the Gaussian spheres [19] or utilize 2D Gaussian disks [15], effectively pulling the Gaussian centers closer to the surfaces of the object. SuGaR [19] incorporates a regularization term that enhances the alignment between Gaussians and surfaces within the scene to improve the accuracy of normal estimation. 2D-GS [15] adopts flattened 2D Gaussians to represent 3D scenes and defines the normal as the direction of the steepest change in density of the 2D Gaussian distribution. GOF [16] utilizes ray-tracing-based volume rendering and establishes a level set by opacities of 3D Gaussians. The surface normals of the Gaussians are approximated using the intersection plane between the ray and the Gaussian, enabling regularization that significantly improves the geometry. These methods solely focus on scenes with consistent light conditions, which often results in the reconstructed mesh exhibiting significant holes and surface inaccuracies. Unlike previous methods, PGSR [25] takes the light conditions into account and incorporates exposure compensation to enhance the accuracy. PGSR flattens the Gaussian into a planar shape and introduces unbiased depth estimation for extracting geometric parameters for surface reconstruction. However, it may result in over-smoothness in the highlight area, and be challenging in reconstructing reflective or mirror surfaces. Therefore, in order to reconstruct meshes under varying light conditions, we propose a method capable of accurate textured mesh extraction in diverse light scenarios.
@@ 3.1 Overview
Figure 2: MGSR is a 2D/3D mutual-boosted framework with two branches: 2D-GS branch (upper) for SR and 3D-GS branch (bottom) for NVS. Each branch is enhanced by our specific designs. Upon receiving inputs from various light conditions, the two branches initially undergo a warm-up stage of initialization for mutual-boosted optimization. In the alternating optimization process under mutual-boosted supervision, the 3D-GS branch is guided by depth maps generated by the 2D-GS branch, while simultaneously, rendered transmitted images without reflections from the 3D-GS branch provide reflection-removed supervision to the 2D-GS branch.
@@
MGSR is a 2D/3D mutual-boosted framework that consists of two branches: improved 2D-GS branch (Sec. 3.3) and 3D-GS branch (Sec. 3.2). Initially, each branch undergoes an independent warm-up phase, after which they engage in alternating optimization through mutual-boosted supervision (Sec. 3.4).
@@
The 3D-GS branch is designed to perform illumination decomposition by introducing reflection-related parameters to disentangle reflections and handle overexposure. However, accurate illumination decomposition relies on reliable geometry, which the 3D-GS branch alone cannot provide. To address this limitation, we introduce a geometry-guided illumination decomposition module, which leverages depth information from the 2D-GS branch to enhance rendering performance under diverse light conditions.
@@
Specifically, the 2D-GS branch aligns 2D Gaussian disks with the surface, capturing initial and relatively accurate geometry information even under significantly varying light conditions. In the subsequent alternating optimization phase, the 3D-GS branch is supervised by depth maps generated by the 2D-GS branch, while reflection-free images rendered by the 3D-GS branch concurrently supervise the 2D-GS branch, facilitating a mutual-boosted alternating optimization. Finally, following to [15], we employ Truncated Signed Distance Fusion (TSDF) [26] to extract the reconstructed textured meshes.
3.2 Illumination decomposition with 3D-GS
@@
3D-GS is constrained in modeling transparent or translucent materials, such as glass. To enhance the comprehensive modeling of scenes under various light conditions, the entire scene is modeled as composed of transmitted and reflected components. Specifically, we retain the original 3D-GS as the transmitted component, while introducing three reflection-related parameters to represent the reflected component: reflected opacity $\theta_{\text{ref}}\in\mathbb{R}$, reflected confidence $\beta\in[0,1]$, and reflected SH $\mathcal{C}_{\text{ref}}\in\mathbb{R}^{k}$, where $k$ refers to the degrees of freedom.
@@
The reflected confidence $\beta$ represents the probability that an individual Gaussian primitive captures the reflected component. When splatting 3D Gaussians to 2D images, $\beta$ is accumulated as described in Eq. 1 to obtain the pixel-wise reflected confidence $W$, and the transmitted color $C_{\text{tran}}$ and reflected color $C_{\text{ref}}$ are calculated with $\alpha$-blending from front to back, as depicted in:
$W=\sum_{i=1}^{K}\beta^{i}{\alpha_{\text{ref}}^{i}}\prod_{j=1}^{k-1}\left(1-\beta^{j}\right),$ (1)
$C_{\text{tran}}=\sum_{i=1}^{K}c_{\text{tran}}^{i}\alpha_{\text{tran}}^{i}\prod_{j=1}^{i-1}\left(1-\alpha_{\text{tran}}^{j}\right),$ (2)
$C_{\text{ref}}=\sum_{i=1}^{K}c_{\text{ref}}^{i}\alpha_{\text{ref}}^{i}\prod_{j=1}^{i-1}\left(1-\alpha_{\text{ref}}^{j}\right),$ (3)
@@
where $i$ denotes the index of the Gaussian sphere, $\alpha_{\text{tran}}$ and $\alpha_{\text{ref}}$ denote to the transmitted opacity $\theta_{\text{tran}}$ and reflected opacity $\theta_{\text{ref}}$ multiplied by the density of the splatted Gaussian at the pixel location, $c_{\text{tran}}^{i}$ and $c_{\text{ref}}^{i}$ represent the view-dependent transmitted and reflected color calculated from transmitted SH $\mathcal{C}_{\text{tran}}$ and reflected SH $\mathcal{C}_{\text{ref}}$. The rendered color $C$ is calculated according to Eq. 4, combining the transmitted and reflected components, weighted by the reflected confidence $W$:
$C=C_{\text{tran}}+W\times C_{\text{ref}}.$ (4)
@@
$\mathcal{L}_{\text{render}}=\lambda_{1}\mathcal{L}_{1}(C,C_{\text{GT}})+(1-\lambda_{1})\mathcal{L}_{\text{D-SSIM}}(C,C_{\text{GT}}),$ (5)
@@
$\mathcal{L}_{\text{3D}}=\mathcal{L}_{\text{render}}+\lambda_{2}\mathcal{L}_{\text{trans-TV}},$ (6)
@@
However, due to the inherent limitation of 3D-GS, specifically the inaccuracy of the depth map, although this 3D branch successfully performs illumination decomposition, it still cannot extract the surface mesh. Therefore, an additional 2D-GS branch is introduced to provide reliable geometry supervision.
3.3 Surface reconstruction with 2D-GS
2D-GS represents the scene with flattened 2D Gaussian primitives, aligning the centers of the Gaussian disks with the surface. The aim of this section is to utilize these flattened Gaussian primitives to obtain initialized depth maps of scenes under various light conditions, which also serve as supervision for the 3D-GS branch and are refined iteratively through the mutual-boosted stage (Sec. 3.4).

### v2

@@ Abstract
Novel view synthesis (NVS) and surface reconstruction (SR) are essential tasks in 3D Gaussian Splatting (3DGS). Despite recent progress, these tasks are often addressed independently, with GS-based rendering methods struggling under diverse light conditions and failing to produce accurate surfaces, while GS-based reconstruction methods frequently compromise rendering quality. This raises a central question: must rendering and reconstruction always involve a trade-off? To address this, we propose MGSR, a 2D/3D Mutual-boosted Gaussian Splatting for Surface Reconstruction that enhances both rendering quality and 3D reconstruction accuracy. MGSR introduces two branches—one based on 2DGS and the other on 3DGS. The 2DGS branch excels in surface reconstruction, providing precise geometry information to the 3DGS branch. Leveraging this geometry, the 3DGS branch employs a geometry-guided illumination decomposition module that captures reflected and transmitted components, enabling realistic rendering under varied light conditions. Using the transmitted component as supervision, the 2DGS branch also achieves high-fidelity surface reconstruction. Throughout the optimization process, the 2DGS and 3DGS branches undergo alternating optimization, providing mutual supervision. Prior to this, each branch completes an independent warm-up phase, with an early stopping strategy implemented to reduce computational costs. We evaluate MGSR on a diverse set of synthetic and real-world datasets, at both object and scene levels, demonstrating strong performance in rendering and surface reconstruction. Code is available at https://github.com/TsingyuanChou/MGSR.
@@ 1 Introduction
3D Gaussian Splatting (3DGS) [9] has recently gained great attention in computer graphics and 3D vision [3]. By representing scenes as collections of 3D Gaussian primitives, 3DGS offers a more flexible and adaptive representation, enabling accurate and efficient rendering and visualization without relying on neural networks. 3DGS effectively resolves the issues of low training efficiency and insufficient geometric accuracy in previous NeRF methods.
Due to its impressive rendering efficiency and accuracy, 3DGS has been widely applied in fields such as SLAM [15], dynamic real-world or large-scale scene modeling [16, 20], reconstruction [18, 11], navigation [28], manipulation [19, 14, 13], 3D generation [2, 29], and 3D human simulation [17], etc.
@@
Figure 1: MGSR achieves strong NVS and SR results compared with methods based on 2DGS [6] and 3DGS [24]. The input consists of multi-view images captured from various camera positions and angles, under significantly varying light conditions. In some views, the object appears well-lit and clearly visible, while in others, it is poorly lit, resulting in shadows, reflections, and inconsistencies across the images.
Nevertheless, achieving high-fidelity surface reconstruction (SR) and improving the realism of novel view synthesis (NVS) under various light conditions are two main challenges in 3DGS.
@@
Specifically, considering that GS-based rendering methods are affected by different light conditions, previous methods [4, 8] proposed to decompose light to enhance the neural rendering in scenes containing reflective surfaces with the introduction of BRDF (Figure 1a). However, despite the effectiveness of illumination decomposition in rendering, these methods are time-consuming and still struggle to achieve meaningful mesh extraction due to inherent limitations in 3DGS. For instance, the centers of Gaussian primitives do not align well with the surface. On the other hand, GS-based SR methods [5, 6, 24] endeavor to constrain the Gaussian primitives close to the surface, inevitably sacrifice rendering quality and are sensitive to light conditions, which may cause surface artifacts on surfaces (Figure 1b). For example, 2DGS [6] utilize 2D Gaussian primitives for 3D scene representation, facilitating accurate and view-consistent geometry modeling. However, 2DGS still cannot effectively model surfaces when ambient lighting changes.
@@
To solve these contradictions, we propose MGSR, a 2D/3D Mutual-boosted Gaussian splatting for Surface Reconstruction that enhances both rendering quality and 3D reconstruction accuracy (Figure 1c). Specifically, two branches are devised based on 2DGS and 3DGS, which are optimized synchronously and mutually enhance each other: the 2DGS branch provides accurate geometric information, with its limited rendering capabilities improved by 3DGS branch, while the 3DGS branch focuses on rendering, with the cost of geometric accuracy compensated by the 2DGS branch. With geometry information, a geometry-guided illumination decomposition module is devised in the 3DGS branch to obtain the reflected and transmitted components, achieving realistic rendering under various light conditions. To achieve this, an additional Spherical Harmonic (SH) is introduced to model reflected colors, along with two other parameters: reflected opacity and reflected confidence, to represent the reflected component accurately. Using the transmitted component as supervision, the 2DGS branch can achieve high-fidelity SR while avoiding the impact of illumination factors on surface estimation. Throughout the mutual-boosted supervision stage, the 2DGS and 3DGS branches engage in alternating optimization for mutual supervision. Prior to alternating optimization, the two modules undergo an independent warm-up stage, and an auto-stop strategy is introduced to reduce unnecessary computational burdens. To the best of our knowledge, MGSR is the first GS-based approach that investigates the simultaneous enhancement of rendering and reconstruction, as well as the first mutual-boosted work on GS involving both 2DGS and 3DGS. MGSR is thoroughly evaluated across a wide range of synthetic and real-world datasets, as well as object- and scene-level datasets, showcasing its superior performance in both rendering and mesh extraction. In summary, our main contributions are as follows:
@@
  • The 2DGS branch aims to provide geometry information to enhance the illumination decomposition of the 3DGS. The decomposed transmitted color will be in turn utilized to supervise the 2DGS branch for improved surface reconstruction, independent of varying light conditions.
  • Leveraging geometric information from the 2DGS branch, we have developed a geometry-guided illumination decomposition module to enhance the accuracy of decomposing reflected and transmitted components, achieving more realistic rendering outcomes.
@@ 2 Related work
Recently, several attempts have also been made at lighting estimation and decomposition on 3DGS. R3DG [4] leverages 3DGS and NEILF [23] to create a ray-tracing and relighting-capable 3DGS representation. GShader [8] presents a simplified shading function for reflective surfaces in 3DGS, where a novel normal estimation is introduced that uses the shortest axis direction of the 3D Gaussian as an approximate normal, eliminating reliance on depth map priors and avoiding the flattening or 2D projection of 3D Gaussians. GS-IR [12] utilizes depth-derivation-based regularization for normal estimation and a baking-based method for modeling indirect lighting, enabling a precise decomposition of material attributes and illumination, thereby significantly improving the photorealism of the rendered outcomes. The prevailing focus in current efforts on lighting estimation and decomposition is predominantly on achieving photorealistic rendering, which consequently hinders the ability to carry out surface mesh extraction effectively.
3DGS driven surface reconstruction. The misalignment of the centroids (means) of 3D Gaussians with the actual surfaces presents a challenge for accurately reconstructing surfaces. A common strategy is to flatten the Gaussian spheres [5] or utilize 2D Gaussian disks [6], effectively pulling the Gaussian centers closer to the surfaces of the object. SuGaR [5] incorporates a regularization term that enhances the alignment between Gaussians and surfaces within the scene to improve the accuracy of normal estimation. 2DGS [6] adopts flattened 2D Gaussians to represent 3D scenes and defines the normal as the direction of the steepest change in density of the 2D Gaussian distribution. GOF [24] utilizes ray-tracing-based volume rendering and establishes a level set by opacities of 3D Gaussians. Surface normals of the Gaussians are approximated using the intersection plane between the ray and the Gaussian, enabling regularization that significantly improves the geometry. These methods solely focus on scenes with consistent light conditions, which often results in the reconstructed mesh exhibiting significant holes and surface inaccuracies. Unlike previous methods, PGSR [1] takes the light conditions into account and incorporates exposure compensation to enhance the accuracy. PGSR flattens the Gaussian into a planar shape and introduces unbiased depth estimation for extracting geometric parameters for surface reconstruction. However, it may result in over-smoothness in the highlight area, and be challenging in reconstructing reflective or mirror surfaces. Therefore, to reconstruct meshes under varying light conditions, we propose a method capable of accurate textured mesh extraction in diverse light scenarios.
@@ 3.1 Overview
Figure 2: MGSR is a 2D/3D mutual-boosted framework with two branches: 2DGS branch (upper) for SR and 3DGS branch (bottom) for NVS. Each branch is enhanced by our specific designs. Upon receiving inputs from various light conditions, the two branches initially undergo a warm-up stage of initialization for mutual-boosted optimization. In the alternating optimization process under mutual-boosted supervision, the 3DGS branch is guided by depth maps generated by the 2DGS branch, while simultaneously, rendered transmitted images without reflections from the 3DGS branch provide reflection-removed supervision to the 2DGS branch.
@@
MGSR is a 2D/3D mutual-boosted framework that consists of two branches: improved 3DGS branch (Section 3.2) and 2DGS branch (Section 3.3). Initially, each branch undergoes an independent warm-up phase, after which they engage in alternating optimization through mutual-boosted supervision (Section 3.4).
@@
The 3DGS branch is designed to perform illumination decomposition by introducing reflection-related parameters to disentangle reflections and handle overexposure. However, accurate illumination decomposition relies on reliable geometry, which the 3DGS branch alone cannot provide. To address this limitation, we introduce a geometry-guided illumination decomposition module, which leverages depth information from the 2DGS branch to enhance rendering performance under diverse light conditions.
@@
Specifically, the 2DGS branch aligns 2D Gaussian disks with the surface, capturing initial and relatively accurate geometry information even under significantly varying light conditions. In the subsequent alternating optimization phase, the 3DGS branch is supervised by depth maps generated by the 2DGS branch, while reflection-free images rendered by the 3DGS branch concurrently supervise the 2DGS branch, facilitating a mutual-boosted alternating optimization. Finally, following to [6], we employ Truncated Signed Distance Fusion (TSDF) [27] to extract the reconstructed textured meshes.
3.2 Illumination decomposition with 3DGS
@@
3DGS is constrained in modeling transparent or translucent materials, such as glass. To enhance the comprehensive modeling of scenes under various light conditions, the entire scene is modeled as composed of transmitted and reflected components. Specifically, we retain the original 3DGS as the transmitted component, while introducing three reflection-related parameters to represent the reflected component: reflected opacity $\theta_{\text{ref}}\in\mathbb{R}$, reflected confidence $\beta\in[0,1]$, and reflected SH $\mathcal{C}_{\text{ref}}\in\mathbb{R}^{k}$, where $k$ refers to the degrees of freedom.
@@
The reflected confidence $\beta$ represents the probability that an individual Gaussian primitive captures the reflected component. When splatting 3D Gaussians to 2D images, $\beta$ is accumulated as described in Equation 1 to obtain the pixel-wise reflected confidence $W$, and the transmitted color $C_{\text{tran}}$ and reflected color $C_{\text{ref}}$ are calculated with $\alpha$-blending from front to back, as depicted in:
$\small W=\sum_{i=1}^{K}\beta^{i}{\alpha_{\text{ref}}^{i}}\prod_{j=1}^{k-1}\left(1-\beta^{j}\right),$ (1)
$\small C_{\text{tran}}=\sum_{i=1}^{K}c_{\text{tran}}^{i}\alpha_{\text{tran}}^{i}\prod_{j=1}^{i-1}\left(1-\alpha_{\text{tran}}^{j}\right),$ (2)
$\small C_{\text{ref}}=\sum_{i=1}^{K}c_{\text{ref}}^{i}\alpha_{\text{ref}}^{i}\prod_{j=1}^{i-1}\left(1-\alpha_{\text{ref}}^{j}\right),$ (3)
@@
where $i$ denotes the index of the Gaussian sphere, $\alpha_{\text{tran}}$ and $\alpha_{\text{ref}}$ denote to the transmitted opacity $\theta_{\text{tran}}$ and reflected opacity $\theta_{\text{ref}}$ multiplied by the density of the splatted Gaussian at the pixel location, $c_{\text{tran}}^{i}$ and $c_{\text{ref}}^{i}$ represent the view-dependent transmitted and reflected color calculated from transmitted SH $\mathcal{C}_{\text{tran}}$ and reflected SH $\mathcal{C}_{\text{ref}}$. The rendered color $C$ is calculated according to Equation 4, combining the transmitted and reflected components, weighted by the reflected confidence $W$:
$\small C=C_{\text{tran}}+W\times C_{\text{ref}}.$ (4)
@@
$\small\mathcal{L}_{\text{render}}=\lambda_{1}\mathcal{L}_{1}(C,C_{\text{GT}})+(1-\lambda_{1})\mathcal{L}_{\text{D-SSIM}}(C,C_{\text{GT}}),$ (5)
@@
$\small\mathcal{L}_{\text{3D}}=\mathcal{L}_{\text{render}}+\lambda_{2}\mathcal{L}_{\text{trans-TV}},$ (6)
@@
However, due to the inherent limitation of 3DGS, specifically the inaccuracy of the depth map, although this 3D branch successfully performs illumination decomposition, it still cannot extract the surface mesh. Therefore, an additional 2DGS branch is introduced to provide reliable geometry supervision.
3.3 Surface reconstruction with 2DGS
2DGS represents the scene with flattened 2D Gaussian primitives, aligning the centers of the Gaussian disks with the surface. The aim of this section is to utilize these flattened Gaussian primitives to obtain initialized depth maps of scenes under various light conditions, which also serve as supervision for the 3DGS branch and are refined iteratively through the mutual-boosted stage (Section 3.4).